# MemoryBench: A Benchmark for Memory and Continual Learning in LLM Systems

### v2

@@ 1 Introduction
Large Language Model (LLM) has been widely considered to be one of the core techniques to achieve Artificial General Intelligence (AGI). By scaling up training data and model parameters, previous studies have shown that LLMs can obtain richer knowledge and stronger abilities in a variety of downstream applications without task-specific fine-tuning, which phenomenon is referred to as the scaling law (Kaplan et al., 2020). Unfortunately, recent experiments show that the scaling law of LLM pre-training hits a wall as we are running out of high-quality data and the gains obtained from further increasing parameters become marginal. While teaching LLMs to conduct online reasoning with more tokens (i.e., the inference scaling law (Wu et al., 2025b)) can further improve the system’s performance, its potential is bounded by the backbone LLMs  (Yang et al., 2025b), and how to train LLMs to obtain such reasoning ability itself is a challenging task that require significant engineering efforts (Yang et al., 2025a). Thus, how to continually improve with LLM systems (LLMsys) and achieve higher level of intelligence remains to be an unsolved question.
@@
In fact, from a broader perspective, the concept of continual learning is not new to human and traditional AI systems. Well-known examples include human brains and search engines. Neuroscience shows that people’s intelligence could continually grow by interacting with the environment even though the number of neurons in their brains is mostly fixed after the age of 25 (Arain et al., 2013). Modern IR systems such as search engines and recommendation systems can continually improve their performance simply by serving and learning from their interactions with online users. By comparing existing LLMsys with human brains and search engines, we can identify two key components behind the ability of continual learning. The first is a mature and stable memory architecture. Human brain has well-developed hippocampus that maintains both context (i.e., short-term memory) and historical (i.e., long-term memory) information, while search engines have well-crafted data and algorithm modules to process dynamic corpus and queries. The second is an effective feedback analysis and utilization mechanism. Human brain can effectively learn knowledge and obtain experience through trial and error, which would then serve as the foundation for them to develop better strategies and solutions for tasks in future; Search engines can utilize different behavior analysis and learning algorithms to understand user’s feedback on search results, which could then be used as features and training data for system optimization (Croft et al., 2010). Therefore, the key to build a LLMsys with continual learning abilities lays in the development of effective memory architectures and feedback utilization (Wu et al., 2025c; Shan et al., 2025; Shi et al., 2025).
@@
Unfortunately, while memory and continual learning have been widely believed to be important and promising research directions, to the best of our knowledge, there isn’t a well-developed benchmark to evaluate the progress of existing algorithms and guide future studies. As discussed in Section 2, existing benchmark on memory for LLMsys mostly focuses on examining LLM system’s ability to handle long context data such as user profiles and conversation histories. They often adopt homogeneous tasks to evaluate LLM systems by simply feeding the context to the memory of LLMsys and ask questions accordingly (Zhang et al., 2025d). More importantly, these benchmarks ignore the dynamic nature of continual learning by evaluating different systems mostly in a static manner (Hu et al., 2025; Maharana et al., 2024; Kim et al., 2025; Zhang et al., 2024; Wu et al., 2025a). They focus on the retrieval of pre-fetched semantic and episodic memory, and do not support the simulation or evaluation of procedural memory built from test-time user feedback. Without feedback, one cannot learn from its or others’ success and failure. This makes these benchmarks not suitable for evaluating the continual learning abilities of LLMsys.
@@
In this paper, we propose a comprehensive benchmark named MemoryBench for the evaluation of memory and continual learning in LLMsys. Inspired by neuroscience (Atkinson & Shiffrin, 1968) and IR (Shen et al., 2005; Agichtein et al., 2006) studies, we categorize the needs of memory based on data characteristics (i.e., declarative or procedural) and build a simulation platform to test the ability of LLMsys in memorizing and learning from user feedback. Our goal is to mimic the process that systems continually learn and improve their performance by interacting with users in online services. To this end, we first collect benchmark datasets in multiple domains, multiple languages, and with multiple tasks that have various lengths of inputs and outputs. Then we split each dataset into separate training and testing sets, and simulate explicit (e.g., a sentence describing whether the output is good or not) and implicit (e.g., a click on the “copy” button) user feedback on the training data using a LLM-as-user paradigm. LLMsys is tested on the testing data to evaluate their abilities in utilizing both the initial input context (i.e., the corpus) and the feedback logs to achieve better task performance. We tested both the state-of-the-art (SOTA) LLM memory systems and naive baselines such as retrieval augmented generation (RAG). Experimental results show that existing memory systems for LLM are not good at utilizing procedural knowledge to improve their performance. They are neither effective nor efficient enough to continually learn from user feedback. More advanced parametric and non-parametric algorithms are still needed in order to enable LLM systems to conduct effective continual learning. For reproducibility, all data processing scripts, user simulators, evaluation pipelines, and tested baselines are open-sourced [^1] https://github.com/LittleDinoC/MemoryBench. We also release simulated user feedback logs for open-sourced LLMs to support off-policy learning experiments [^2] https://huggingface.co/datasets/THUIR/MemoryBench. We hope this could pave the way for future studies on memory and continual learning for LLMsys.
@@ 2.1 Preliminary and Taxonomy
The concept of memory in intelligent systems has been extensively studied in the field of both neural science, social science, and computer science. For example, previous studies have developed different taxonomies for memory based on what they store (e.g., facts, events, feelings, etc.) and how they are managed (e.g., short-term, long-term, etc.) (Atkinson & Shiffrin, 1968). In this paper, our goal is to evaluate different LLMsys with or without memory but not developing new architectures or algorithms. Thus, inspired by previous studies (Cohen & Bacdayan, 1994; Tulving & Markowitsch, 1998), we analyze memory based on the knowledge they store and categorize them into two types: Declarative Memory and Procedural Memory. Declarative memory refers to factual knowledge independent to tasks. Depending on whether the knowledge is related to users, we further categorize it into Semantic Memory and Episodic Memory. We refer to semantic memory as data and knowledge independent to users, such as textbook, wikipedia, etc. We refer to episodic memory as memory that is user dependent, such as conversational history, personal profiles, etc. In contrast to declarative memory, procedural memory focuses on non-factual knowledge related to the execution of tasks, such as the workflow of a task, the reward of a solution, etc. User feedback and behavior logs can be considered as important resources for the procedural memory of LLMsys.
In contrast to memory, the taxonomy of user feedback is not well explored in existing LLM literature. While the word “feedback” is not new for LLM, most existing works use it as a general term for training and testing reward rather than user behavior signals collected in LLM services (Ouyang et al., 2022). Inspired by user studies in IR (Shen & Zhai, 2005; White et al., 2006), we categorize user feedback for LLMsys into two types: Explicit feedback and Implicit Feedback. Explicit feedback refers to user behaviors that directly reflect the quality of the LLMsys’s output. For example, a user may conduct multi-turn conversations with a LLMsys by directly telling it what’s wrong with its output and how to fix it, which we refer to as verbose feedback. A user may also tell the system whether its response is satisfying by simply clicking the “like” or “dislike” button widely provided in chatbot user interfaces, which we refer to as action feedback. In contrast, implicit feedback denotes user behavior actions generated during the use of LLM services, but not for the purpose of judging the quality of LLM responses. Examples include clicking the “copy” button, closing a session after seeing the response, or starting a new session with a similar but refined prompt. These behaviors could be useful for optimizing the system, but may contain noise and need significant analysis in order to extract reliable reward signals. Based on this taxonomy, we analyze and categorize existing benchmarks on LLM agent and memory in Table 1. To the best of our knowledge, MemoryBench is the first benchmark that provides all types of memory and feedback data. Also, previous benchmarks mainly focus on testing LLMsys on answering factual questions where the answers can be extracted from the given context (i.e., the corpus), which are essentially reading comprehension tasks with long input and short output. In contrast, in MemoryBench, we include diverse tasks from multiple domains that have various lengths of input and output (details in A.1). These make MemoryBench a unique resource to evaluate the continual learning ability of LLMsys.
Table 1: Benchmarks Categorization and Comparison. Dec., Pro., Sem., Epi., Exp., Imp., Ver., Act. are abbreviations for Declarative, Procedural, Semantic, Episodic, Explicit, Implicit, Verbose, and Action. LiSo is the acronym of Long-input-Short-output, etc.
@@
Hu et al. (2025) LiSo $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 2k
Wu et al. (2025a) LiSo $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 0.5k
Kim et al. (2025) LiSo $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 3121k
Maharana et al. (2024) LiSo $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 7k
Du et al. (2024) LiSo $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 8k
Zhang et al. (2024) LiSo $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 3k
@@ 2.2 Architecture
Performance Monitor is designed to monitor the performance of LLMsys on the test data. It examines whether LLMsys can improve their task performance through interacting with the simulated users. Specifically, we adopt the native metrics used by the original datasets for evaluation. Because different tasks may involve multiple and diverse evaluation metrics, we implement a LLM-as-judge paradigm (Li et al., 2024b) to merge multiple metrics into one performance score per task case. We aggregate multiple task scores by first conducting a min-max normalization or computing the z-score of all tested LLMsys and test cases within each dataset, and then averaging the performance score of a specific LLMsys on test cases from all datasets to compute the final score for this LLMsys. Note that the performance monitor evaluates LLMsys only on the test data, which are different from the training data fed by the task provider to the user simulator.
@@ 2.3 Data Preparation
To analyze LLMsys’s ability in utilizing declarative and procedural memory in heterogeneous tasks, we collect 11 public datasets across three domains, four types of task formats, and two languages. Specifically, they are Locomo (Maharana et al., 2024), DialSim (Kim et al., 2025), LexEval (Li et al., 2024a), JuDGE (Su et al., 2025), IdeaBench (Guo et al., 2025b), LimitGen-Syn (Xu et al., 2025b), WritingPrompts (Fan et al., 2018), HelloBench (Que et al., 2024), WritingBench (Wu et al., 2025d), NF-Cats (Bolotova et al., 2022), and SciTechNews (Cardenas et al., 2023). As shown in Table 2, these datasets cover open-domain, legal, and academic data, and their tasks exhibit wide variations in input and output formats and lengths. We believe that these datasets can closely reflect many daily usage scenarios for LLMsys, which are suitable for the generation and collection of user feedback logs. For datasets that span multiple tasks and domains (e.g., WritingBench), we further divide them into subsets based on our categorization taxonomy. For all datasets, we unify their representations and format each task case as $(q,v,c)$, where $q$ denotes the input query or instruction submitted by the user, $v$ represents the metadata needed for evaluation (e.g., datasets such as JuDGE and WritingBench rely on either golden answers or multiple hand-crafted criteria to evaluate LLMsys’s responses), and $c$ is the task context, i.e., the corpus, needed to answer the query. Here, $c$ is optional as it depends on the task and dataset. For example, Locomo and DialSim require LLMsys to answer questions based on user’s previous dialogs, and these dialogs are considered as $c$ for the task. Also, depending on the tasks, $c$ could be user-dependent or user-independent. For simplicity, we do not differentiate it explicitly and let the LLMsys decide how to process the task context. The design of $v$ is closely related to the main test metrics used by each dataset. For datasets that have multiple metrics (e.g., IdeaBench, JuDGE, and SciTechNews), we adopt a LLM-as-judge paradigm and merge those metrics into one. Please refer to A.1 for more details.
@@ 2.4 Feedback Simulation
  • Action Feedback: The simulator predicts a user satisfaction score for the LLMsys’s response and generates explicit and implicit actions such as “like”, “dislike”, or “copy” based on mapping functions. The functions are task-dependent and programmable so that the generated actions could follow diverse user behavior patterns. Specifically, we design the mapping functions following the observations obtained from recent studies on user behaviors on LLM services (Shuster et al., 2022; Chatterji et al., 2025). More details can be found in A.2.
@@ 3.1 Baseline Categorization
To demonstrate how to use MemoryBench, we implemented and tested simple baselines as well as a couple of SOTA memory-based LLMsys, which are naive retrieval augmented generation (RAG) with BM25 and Qwen3-Embedding-0.6B (Zhang et al., 2025c) as the retrievers, A-Mem (Xu et al., 2025a), Mem0 (Chhikara et al., 2025), and MemoryOS (Kang et al., 2025). We use the same backbone LLM (i.e., Qwen3-8B, but we have also tested Qwen3-32B as shown in Table 10 in A.3) for all baselines and refer to the naive method that directly answers each query with the backbone LLM as Vanilla. For memory-based LLMsys, we directly adopt their original code and design from their papers to manage input corpus and feedback logs. For RAG methods, we store the corpus directly and tested two methods to process the feedback logs: forming a single document with the whole dialog in each session, or storing each message in a session as separate documents. We refer to the baselines using BM25 and Qwen3-Embedding-0.6B with the Session documents and Message documents as BM25-S, BM25-M, Embed-S, and Embed-M, respectively. For action-based feedback such as “like” and “copy”, we also implemented a supervised fine-tuning baselines (SFT) and reported its performance in A.3.9. When the input context of a case is longer than the input limit of a baseline, we simply cut the context to fit. More details can be found in A.3.

### v4

@@ 1 Introduction
Large Language Model (LLM) has been widely considered to be one of the core techniques to achieve Artificial General Intelligence (AGI). By scaling up training data and model parameters, previous studies have shown that LLMs can obtain richer knowledge and stronger abilities in a variety of downstream applications without task-specific fine-tuning. This phenomenon is referred to as the scaling law (kaplan2020scaling). Unfortunately, recent experiments show that the scaling law of LLM pre-training hits a wall as we are running out of high-quality data and the gains obtained from further increasing parameters become marginal. While teaching LLMs to conduct online reasoning with more tokens (i.e., the inference scaling law (wu2025inferencescalinglawsempirical)) can further improve the system’s performance, its potential is bounded by the backbone LLMs  (yang2025probabilisticinferencescalingtheory), and how to train LLMs to obtain such reasoning ability itself is a challenging task that require significant engineering efforts (yang2025qwen3). Thus, how to continually improve LLM systems (LLMsys) and achieve higher level of intelligence remains to be an unsolved question.
@@
In fact, from a broader perspective, the concept of continual learning is not new to human and traditional AI systems. Well-known examples include human brains and search engines. Neuroscience shows that people’s intelligence could continually grow by interacting with the environment even though the number of neurons in their brains is mostly fixed after the age of 25 (arain2013maturation). Modern IR systems such as search engines and recommendation systems can continually improve their performance simply by serving and learning from their interactions with online users. By comparing existing LLMsys with human brains and search engines, we can identify two key components behind the ability of continual learning. The first is a mature and stable memory architecture. Human brain has well-developed hippocampus that maintains both context (i.e., short-term memory) and historical (i.e., long-term memory) information, while search engines have well-crafted data and algorithm modules to process dynamic corpus and queries. The second is an effective feedback analysis and utilization mechanism. Human brain can effectively learn knowledge and obtain experience through trial and error, which would then serve as the foundation for them to develop better strategies and solutions for tasks in the future; Search engines can utilize different behavior analysis and learning algorithms to understand user’s feedback on search results, which could then be used as features and training data for system optimization (croft2010search). Therefore, the key to build a LLMsys with continual learning abilities lays in the development of effective memory architectures and feedback utilization (wu2025humanmemoryaimemory; shan2025cognitivememorylargelanguage; shi2025continuallearning).
@@
Unfortunately, while memory and continual learning have been widely believed to be important and promising research directions, to the best of our knowledge, there is no well-developed benchmark to evaluate the progress of existing algorithms and guide future studies. As discussed in Section 2, existing benchmark on memory for LLMsys mostly focuses on examining LLM system’s ability to handle long context data such as user profiles and conversation histories. They often adopt homogeneous tasks to evaluate LLM systems by simply feeding the context to the memory of LLMsys and ask questions accordingly (ZhangSurvey). More importantly, these benchmarks ignore the dynamic nature of continual learning by evaluating different systems mostly in a static manner (hu2025evaluatingmemoryllmagents; maharana2024evaluatinglongtermconversationalmemory; kim2025dialsimrealtimesimulatorevaluating; zhang2024memsimbayesiansimulatorevaluating; wu2025longmemeval; tan2025membench). They focus on the retrieval of pre-fetched semantic and episodic memory, and do not support the simulation or evaluation of procedural memory built from test-time user feedback. Without feedback, one cannot learn from its or others’ success and failure. This makes these benchmarks not suitable for evaluating the continual learning abilities of LLMsys.
@@
In this paper, we propose a comprehensive benchmark named MemoryBench for the evaluation of memory and continual learning in LLMsys. Inspired by neuroscience (atkinson1968human) and IR (shen2005implicit; Eugene2026websearch) studies, we categorize the needs of memory based on data characteristics (i.e., declarative or procedural) and build a simulation platform to test the ability of LLMsys in memorizing and learning from user feedback. Our goal is to mimic the process that systems continually learn and improve their performance by interacting with users in online services. To this end, we first collect benchmark datasets in multiple domains, multiple languages, and with multiple tasks that have various lengths of inputs and outputs. Then we split each dataset into separate training and testing sets, and simulate explicit (e.g., a sentence describing whether the output is good or not) and implicit (e.g., a click on the “copy” button) user feedback on the training data using a LLM-as-user paradigm. LLMsys is tested on the testing data to evaluate their abilities in utilizing both the initial input context (i.e., the corpus) and the feedback logs to achieve better task performance. We tested both the state-of-the-art (SOTA) LLM memory systems and naive baselines such as retrieval augmented generation (RAG). Experimental results show that existing memory systems for LLM are not good at utilizing procedural knowledge to improve their performance. They are neither effective nor efficient enough to continually learn from user feedback. More advanced parametric and non-parametric algorithms are still needed in order to enable LLM systems to conduct effective continual learning. For reproducibility, all data processing scripts, user simulators, evaluation pipelines, and tested baselines are open-sourced [^1] https://github.com/LittleDinoC/MemoryBench. We also release simulated user feedback logs for open-sourced LLMs to support off-policy learning experiments [^2] https://huggingface.co/datasets/THUIR/MemoryBench. We hope this could pave the way for future studies on memory and continual learning for LLMsys.
@@ 2.1 Preliminary and Taxonomy
Table 1: Benchmarks Categorization and Comparison. Dec., Pro., Sem., Epi., Exp., Imp., Ver., Act. are abbreviations for Declarative, Procedural, Semantic, Episodic, Explicit, Implicit, Verbose, and Action. LiSo is the acronym of Long-input-Short-output, etc. #Case is the number of task cases in the corresponding dataset.
@@
hu2025evaluatingmemoryllmagents LiSo $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 2k
wu2025longmemeval LiSo $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 0.5k
kim2025dialsimrealtimesimulatorevaluating LiSo $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 3121k
maharana2024evaluatinglongtermconversationalmemory LiSo $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 7k
du2024perltqapersonallongtermmemory LiSo $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 8k
zhang2024memsimbayesiansimulatorevaluating LiSo $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 3k
tan2025membench LiSo $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 53k
@@
The concept of memory in intelligent systems has been extensively studied in the field of both neural science, social science, and computer science. For example, previous studies have developed different taxonomies for memory based on what they store (e.g., facts, events, feelings, etc.) and how they are managed (e.g., short-term, long-term, etc.) (atkinson1968human). In this paper, our goal is to evaluate different LLMsys with or without memory but not developing new architectures or algorithms. Thus, inspired by previous studies (cohen1994organizational; tulving1998episodic), we analyze memory based on the knowledge they store and categorize them into two types: Declarative Memory and Procedural Memory. Declarative memory refers to factual knowledge independent to tasks. Depending on whether the knowledge is related to users, we further categorize it into Semantic Memory and Episodic Memory. We refer to semantic memory as data and knowledge independent to users, such as textbook, wikipedia, or other factual information collected from the world, etc. We refer to episodic memory as memory that is user dependent, such as conversational history, personal profiles, or other user-specific information, etc. In contrast to declarative memory, procedural memory focuses on non-factual knowledge related to the execution of tasks, such as the workflow of a task, the reward of a solution, etc. For instance, user feedback and behavior logs that contain important information about how the system could improve its performance in the future can be considered as important resources for the procedural memory of LLMsys. An illustrative example is shown in Figure 1 to better show the relationships and differences between different types of memory information.
Figure 1: An illustrative example of different types of memory and user feedback.
In contrast to memory, the taxonomy of user feedback is not well explored in existing LLM literature. While the word “feedback” is not new for LLM, most existing works use it as a general term for training and testing reward rather than user behavior signals collected in LLM services (ouyang2022training). Inspired by user studies in IR (shen2005feedback; WHITE2006166), we categorize user feedback for LLMsys into two types: Explicit feedback and Implicit Feedback. Explicit feedback refers to user behaviors that directly reflect the quality of the LLMsys’s output. For example, a user may conduct multi-turn conversations with a LLMsys by directly telling it what’s wrong with its output and how to fix it, which we refer to as verbose feedback. A user may also tell the system whether its response is satisfying by simply clicking the “like” or “dislike” button widely provided in chatbot user interfaces, which we refer to as action feedback. An example can be found in Figure 1. In contrast, implicit feedback denotes user behavior actions generated during the use of LLM services, but not for the purpose of judging the quality of LLM responses. Examples include clicking the “copy” button, closing a session after seeing the response, or starting a new session with a similar but refined prompt. These behaviors could be useful for optimizing the system, but may contain noise and need significant analysis in order to extract reliable reward signals. Based on this taxonomy, we analyze and categorize existing benchmarks on LLM agent and memory in Table 1. To the best of our knowledge, MemoryBench is the first benchmark that provides all types of memory and feedback data. Also, previous benchmarks mainly focus on testing LLMsys on answering factual questions where the answers can be extracted from the given context (i.e., the corpus), which are essentially reading comprehension tasks with long input and short output. In contrast, in MemoryBench, we include diverse tasks from multiple domains that have various lengths of input and output (details in A.1). These make MemoryBench a unique resource to evaluate the continual learning ability of LLMsys.
@@ 2.2 Architecture
Performance Monitor is designed to monitor the performance of LLMsys on the test data. It examines whether LLMsys can improve their task performance through interacting with the simulated users. Specifically, we adopt the native metrics used by the original datasets for evaluation. Because different tasks may involve multiple and diverse evaluation metrics, we implement a LLM-as-judge paradigm (li2024llmsasjudgescomprehensivesurveyllmbased) to merge multiple metrics into one performance score per task case. We aggregate multiple task scores by first conducting a min-max normalization or computing the z-score of all tested LLMsys and test cases within each dataset, and then averaging the performance score of a specific LLMsys on test cases from all datasets to compute the final score for this LLMsys. Note that the performance monitor evaluates LLMsys only on the test data, which are different from the training data fed by the task provider to the user simulator.
@@ 2.3 Data Preparation
To analyze LLMsys’s ability in utilizing declarative and procedural memory in heterogeneous tasks, we collect 11 public datasets across three domains, four types of task formats, and two languages. Specifically, they are Locomo (maharana2024evaluatinglongtermconversationalmemory), DialSim (kim2025dialsimrealtimesimulatorevaluating), LexEval (NEURIPS2024_2cb40fc0-lexeval), JuDGE (judge2025), IdeaBench (ideabench), LimitGen-Syn (xu-etal-2025-llms-identify-limitgen), WritingPrompts (fan-etal-2018-hierarchical-writingprompts), HelloBench (que2024hellobenchevaluatinglongtext), WritingBench (wu2025writingbench), NF-Cats (nfcats), and SciTechNews (cardenas-etal-2023-dont-scitechnews). As shown in Table 2, these datasets cover open-domain, legal, and academic data, and their tasks exhibit wide variations in input and output formats and lengths. We believe that these datasets can closely reflect many daily usage scenarios for LLMsys, which are suitable for the generation and collection of user feedback logs. For datasets that span multiple tasks and domains (e.g., WritingBench), we further divide them into subsets based on our categorization taxonomy. For all datasets, we unify their representations and format each task case as $(q,v,c)$, where $q$ denotes the input query or instruction submitted by the user, $v$ represents the metadata needed for evaluation (e.g., datasets such as JuDGE and WritingBench rely on either golden answers or multiple hand-crafted criteria to evaluate LLMsys’s responses), and $c$ is the task context, i.e., the corpus, needed to answer the query. Here, $c$ is optional as it depends on the task and dataset. For example, Locomo and DialSim require LLMsys to answer questions based on user’s previous dialogs, and these dialogs are considered as $c$ for the task. Also, depending on the tasks, $c$ could be user-dependent or user-independent. For simplicity, we do not differentiate it explicitly and let the LLMsys decide how to process the task context. The design of $v$ is closely related to the main test metrics used by each dataset. For datasets that have multiple metrics (e.g., IdeaBench, JuDGE, and SciTechNews), we adopt a LLM-as-judge paradigm and merge those metrics into one. Please refer to A.1 for more details.
@@ 2.4 Feedback Simulation
  • Action Feedback: The simulator predicts a user satisfaction score for the LLMsys’s response and generates explicit and implicit actions such as “like”, “dislike”, or “copy” based on mapping functions. The functions are task-dependent and programmable so that the generated actions could follow diverse user behavior patterns. Specifically, we design the mapping functions following the observations obtained from recent studies on user behaviors on LLM services (shuster2022blenderbot; chatterji2025people). More details can be found in A.2.
To verify the quality of our LLM-as-user paradigm in terms of generating high-quality verbose feedback, we have conducted human annotation experiments to compare human-written feedback with those generated in MemoryBench, and results show that it is difficult for human annotators to distinguish human feedback with our simulated feedback. More details can be found in A.2.5.
@@ 3.1 Baseline Categorization
To demonstrate how to use MemoryBench, we implemented and tested simple baselines as well as a couple of SOTA memory-based LLMsys, which are naive retrieval augmented generation (RAG) with BM25 and Qwen3-Embedding-0.6B (zhang2025qwen3) as the retrievers, A-Mem (xu2025mem), Mem0 (chhikara2025mem0), and MemoryOS (kang2025memory). We use the same backbone LLM (i.e., Qwen3-8B, but we have also tested Qwen3-32B and Mistral3.2-24B as shown in Table 15 and Table 22 in A.3) for all baselines and refer to the naive method that directly answers each query with the backbone LLM as Vanilla. For memory-based LLMsys, we directly adopt their original code and design from their papers to manage input corpus and feedback logs. For RAG methods, we store the corpus directly and tested two methods to process the feedback logs: forming a single document with the whole dialog in each session, or storing each message in a session as separate documents. We refer to the baselines using BM25 and Qwen3-Embedding-0.6B with the Session documents and Message documents as BM25-S, BM25-M, Embed-S, and Embed-M, respectively. For action-based feedback such as “like” and “copy”, we also implemented a supervised fine-tuning baselines (SFT) and reported its performance in A.3.9. When the input context of a case is longer than the input limit of a baseline, we simply cut the context to fit. More details can be found in A.3.